### v1

# Workflow as a Service Broker in Cloud Environment: A Systematic Literature Review

Abstract.
Cloud computing has emerged as a promising platform for running scientific workflows across various domains. Scientists can take advantage of different cloud service models, such as serverful or serverless, to execute workflows based on their specific requirements, along with diverse pricing models like on-demand, reserved, or spot instances to reduce execution costs. However, the challenge of selecting appropriate resources and pricing models, coupled with the orchestration and scheduling of workflow tasks, creates significant complexity for users. To mitigate this burden, Workflow as a Service (WaaS) brokers have been introduced to facilitate workflow execution. In recent years, numerous studies have been published, either directly or indirectly related to this research area, highlighting the need for a comprehensive and systematic review of WaaS brokers to identify key trends and challenges in this field. In this paper, we conduct a Systematic Literature Review (SLR) on WaaS brokers within cloud environments. The SLR employs a thorough 3-tier strategy (database search, backward snowballing, and forward snowballing) to answer five research questions. A total of 74 high-quality articles, published in 43 prestigious venues, are analyzed to derive a taxonomy based on the architecture of WaaS brokers. The articles are classified and surveyed according to this taxonomy, and future research directions for the design and implementation of WaaS brokers are explored. This study provides valuable insights for researchers and developers, helping them identify major trends and issues in the field of WaaS brokers.
Workflow as a Service, cloud broker, multiple workflow scheduling, systematic literature review
ccs: Computer systems organization Cloud computingccs: Theory of computation Scheduling algorithmsccs: Computing methodologies Distributed computing methodologiesccs: Computer systems organization Architectures Distributed architectures Cloud computing
1. Introduction
Cloud computing has emerged as one of the leading technologies, transforming industries and reshaping the world. It allows customers to access and rent the resources they need on demand, paying on a pay-as-you-go basis instead of purchasing resources outright (Josep et al., 2010). Both industry and academia have embraced this new computational paradigm to meet their resource requirements (Versluis and Iosup, 2021). In particular, new paradigms in Artificial Intelligence, such as deep learning and large language models, can leverage large pools of resources, such as GPUs, to address their computational demands. Furthermore, leading cloud providers, such as Amazon, are striving to introduce new services to cater to evolving user needs and attract new customers. As a result, alongside traditional services like Infrastructure as a Service (IaaS), which provides Virtual Machines (VMs), we now have services such as Container as a Service (CaaS) and Function as a Service (FaaS), or serverless computing (Shafiei et al., 2022), which offer higher-level services beyond raw VMs.
Nowadays, many applications running on cluster and cloud resources take the form of workflows (Versluis and Iosup, 2021). The concept of workflows, originally rooted in the manufacturing and business process domains, has evolved into the scientific community, where it is now referred to as scientific workflows (Hilman et al., 2020). A scientific workflow describes a process aimed at achieving a scientific objective, typically expressed in terms of tasks and their dependencies. Scientific workflows are often represented as a Directed Acyclic Graph (DAG), where each vertex corresponds to a task, and each edge denotes a data or control dependency between two tasks. Researchers from various scientific domains commonly define their applications as scientific workflows (Juve et al., 2013; Callaghan et al., 2024).
Cloud computing provides an attractive environment for executing scientific workflows. Scientists can leverage various cloud services, such as IaaS, CaaS, or FaaS, alongside different pricing models, including on-demand, spot, or reserved, based on their specific needs and requirements, to execute workflows efficiently and cost-effectively. However, selecting appropriate resources and pricing models, and more critically, orchestrating and scheduling workflow tasks on these resources, poses a significant challenge for users. This is where cloud brokers come into play. A broker plays a vital role in addressing these challenges by mediating between cloud providers and customers to identify the most suitable services while considering user preferences (Li et al., 2022b). Additionally, brokers can assist users in avoiding the vendor lock-in problem, which refers to the dependence on a specific cloud provider’s services, potentially leading to monopolistic scenarios.
Workflow as a Service (WaaS) is a type of cloud broker that accepts users’ workflows along with their specified Quality of Service (QoS) requirements, such as deadline and budget, and executes these workflows by provisioning and allocating appropriate resources from cloud provider(s). While the broker seeks to optimize its own benefits, it also aims to provide sufficient incentives for users to delegate their workflows, primarily by reducing the cost of execution. To achieve this, the broker operates through two key components: the resource provisioner and the task scheduler. The resource provisioner is tasked with selecting suitable resources from various service models (IaaS, SaaS, or FaaS) and pricing models (on-demand, spot, and reserved). For example, to minimize costs, the provisioner might rent a large VM and use containerization to share it among tasks from multiple workflows. Additionally, while most users rely on costly on-demand resources, the provisioner can opt for cheaper alternatives, such as spot instances, by managing their unreliability, or reserved instances for workflows that run over extended periods. The scheduler is responsible for mapping the tasks of individual workflows onto the provisioned resources. By running multiple workflows concurrently, the scheduler can optimize resource utilization more effectively than scheduling each workflow independently. Consequently, the scheduling challenge in WaaS brokers is often referred to as the Multiple Workflow (or Multi-Workflow) Scheduling Problem (MWSP).
A comprehensive and systematic review of research on WaaS brokers is essential for identifying major trends and issues in this field. Such a review can serve as a valuable guideline for researchers and enthusiasts, providing a deeper understanding of the progress made so far and the challenges that remain to be addressed. To ensure proper coverage of research studies in the domain of WaaS brokers, a systematic methodology is critical for identifying all relevant search spaces and studies (Petersen et al., 2015; Kitchenham et al., 2009). To the best of our knowledge, no prior Systematic Literature Review (SLR) has been conducted specifically on WaaS brokers. While a few studies address the broader domain of cloud brokerage (Li et al., 2022b; Khorasani et al., 2024), they do not specifically examine WaaS brokers. However, given that the scheduler is a key component of a WaaS broker, surveys on workflow scheduling in cloud environments are considered related works. Several studies have surveyed cloud workflow scheduling in general (Versluis and Iosup, 2021; Adhikari et al., 2019; Rodriguez and Buyya, 2017), but while these articles touch on the Multiple Workflow Scheduling Problem (MWSP), they do not comprehensively address it. Recently, Saeeidzade and Ashtiani (Saeedizade and Ashtiani, 2024) published a survey on workflow scheduling in cloud environments, which includes a subsection on scheduling in WaaS brokers. However, this survey does not comprehensively explore the domain. The most relevant work to this SLR is by Hilman et al. (Hilman et al., 2020), which provides a detailed survey of MWSP in multi-tenant distributed systems such as clusters, grids, and clouds. While this survey is comprehensive, it has several limitations. First, its taxonomy encompasses all distributed systems, including aspects irrelevant to the cloud environment. Second, it was published in 2019 and reviews only 31 articles, whereas numerous new studies have emerged in the past five years, particularly in the field of WaaS brokers (as shown in Section 3). This highlights the need for an updated survey. Finally, new technologies, such as serverless computing, have gained significant attention since the publication of Hilman et al.’s survey, further underscoring the necessity for a new review focused on recent developments in this area.
We have conducted a systematic literature review (SLR) (Kitchenham et al., 2009) in the field of WaaS brokers within the cloud environment. To identify studies related to this area, we designed queries based on key terms relevant to the field. A three-tier search strategy was employed, consisting of database search, backward snowballing, and forward snowballing (see Section 3). By integrating these three search strategies, 74 papers from 43 different venues (journals and conferences) were selected based on specific inclusion and exclusion criteria. These papers were then analyzed to answer 5 research questions, serving as a valuable resource for research teams and developers in the field. All included papers and the extracted information for answering the research questions are available in an Excel supplementary file, accessible at https://github.com/SaeidAbrishami/WaaSBroker.
The remainder of this article is organized as follows: Section 2 provides a detailed introduction to the architecture of WaaS brokers. Section 3 elaborates on the research methodology and provides a brief report of the review. Section 4 proposes a taxonomy for this field and surveys existing solutions based on this taxonomy. The future directions of WaaS brokers are outlined in Section 5, and Section 6 concludes the article.
2. The architecture of WaaS broker
Scientific workflows are a common method for automating and executing complex scientific applications across various fields (Juve et al., 2013; Callaghan et al., 2024). These sophisticated applications typically require execution on resource-rich computing environments, such as distributed infrastructures. Cloud computing offers an ideal infrastructure for running these workflows by providing various services, including IaaS, CaaS, and FaaS, with different pricing models like on-demand, spot, and reserved. However, selecting the appropriate services and resources, along with the correct pricing model, poses a significant challenge for scientists as users. Additionally, efficiently utilizing resources through optimal scheduling of workflow tasks is another challenge for users. Workflow as a Service is an emerging paradigm that allows users (scientists) to execute their scientific workflows without worrying about resource management and task scheduling. WaaS is a type of broker that receives scientific workflows along with their required QoS, then manages the provisioning of necessary resources from cloud providers and schedules workflow tasks on these resources.
Figure 1. The architecture of a Workflow as a Service broker.
Figure 1 illustrates the architecture of a typical WaaS broker. As shown, there are three main actors in the system: users, cloud provider(s), and the WaaS broker. The process begins when users send one or more workflows to the WaaS broker. A workflow can be represented using a DAG, where each node corresponds to a task, and each edge represents a data or control dependency between two tasks. This dependency indicates that a child task can only begin execution once its parent task(s) has completed and the child task has received the necessary output data. Additionally, users can often specify QoS parameters, such as deadlines and maximum budgets, for executing the workflow through a Service Level Agreement (SLA). In some cases, the broker may agree to a (monetary) penalty if the SLA is violated.
The cloud provider offers the necessary services and resources to execute users’ workflows. Currently, three well-known services are provided: IaaS, CaaS, and FaaS. The traditional IaaS service provides VMs to users, with various hardware and software specifications. CaaS, a newer service, offers lightweight virtualized environments known as containers. Containers have become popular for managing applications in the cloud (Pahl et al., 2019). By using a container, a virtual environment is created with separate CPU, memory, block I/O, and network space, allowing multiple tasks to run in parallel on a single VM. These services are typically provided in two models. For example, Amazon Elastic Container Service (ECS) allows users to fully manage containers on their VMs leased through AWS EC2. However, ECS also offers a service called AWS Fargate, which enables users to run containers without worrying about where or on which VM they will be executed; Amazon itself manages the deployment and operation of these containers. Both types of containers have been employed by researchers. Finally, the most recent service provided by cloud providers is serverless computing, typically represented by the FaaS service (along with Backend as a Service (BaaS)). FaaS allows developers to write stateless, event-driven serverless functions, enabling them to focus on the logic of serverless applications without worrying about the resources (VMs or containers) on which the application will run; the provider manages these resources (Shafiei et al., 2022). AWS Lambda is the most well-known model of this service. Shared storage services, such as Amazon S3 and Amazon Elastic File System (EFS), are also essential for storing intermediate data between tasks.
Another important aspect of the services of the cloud provider is the pricing model. On-demand is the oldest and most well-known pricing model: the cloud provider offers various services at pre-determined prices, and the user can lease resources on a pay-as-you-go basis. This means the user launches an instance of a resource whenever needed and terminates it when the task is complete, paying only for the time the resource was used, without any long-term commitment. Historically, users were charged by the hour, but now, major providers such as Amazon charge users per second. In contrast, the reserved pricing model allows the user to lease resources for an extended period (typically one year or more) at a significant discount compared to on-demand resources (around 40% off in Amazon). Finally, the spot pricing model enables users to access the cloud provider’s excess computing capacity at a substantial discount (up to 80% in Amazon). However, these resources are highly unreliable, as the cloud provider can reclaim them with as little as two minutes’ notice. Therefore, the user must plan accordingly to manage this uncertainty (Lin et al., 2022). The final pricing model pertains to the FaaS model, in which the user is typically billed based on the number of function invocations, the time it takes for execution, and the computational resources allocated to the function instance, particularly memory.
Figure 1 shows that the WaaS broker consists of four components: negotiation and pricing, task scheduler, resource provisioner, and monitoring and fault tolerance. The negotiation and pricing component interacts with the user. It receives the user’s workflow along with its required QoS and determines if it can meet the requirements. If the answer is yes, it specifies the price for running the workflow. It is reasonable for the price to be lower than what the user would pay by directly leasing resources from the provider. However, the WaaS broker must also generate a profit. This can be achieved by efficiently using leased resources, sharing them between different workflows, and utilizing pricing models such as reserved instances to reduce the cost of workflow execution.
The task scheduler is responsible for mapping and arranging the workflow’s tasks onto the leased resources, optimizing certain criteria (such as makespan or execution cost) while meeting the workflow’s QoS requirements. To make efficient use of the leased resources, the scheduler should consider all tasks from different workflows simultaneously, rather than scheduling each workflow’s tasks individually. Therefore, the scheduler falls under the domain of Multi-Workflow Scheduling Problem (MWSP). Since the scheduler is the core component of the broker, as discussed in Section 3.2, these terms have been used as search strings to identify related articles. However, we have focused on articles that address the MWSP on leased resources (VMs or containers) from the client-side (or broker-side) and excluded papers that explore this problem from the provider-side, such as those that involve scheduling tasks on physical machines, such as (Wang et al., 2023, 2024; Fan et al., 2023a).
The resource provisioner is responsible for leasing appropriate resources based on the needs of the task scheduler. This component determines when, how many, and which type (in terms of computational power, such as the number of cores and memory) of resources should be requested from the provider, and when they should be released to avoid wasting money. It is important to note that the provisioner is not relevant in FaaS platforms, as resource management is handled by the provider itself. The final component is the monitoring and fault tolerance module, which is responsible for overseeing the correct execution of tasks and the status of the leased resources. Additionally, in the event of a failure, this component must take necessary actions to recover from the issue, such as re-executing the failed tasks. This is especially crucial in the spot pricing model, where resources have low reliability and can be reclaimed by the provider at short notice.
3. Research methodology
We have conducted an SLR on WaaS broker according to the guidelines of (Petersen et al., 2015; Kitchenham et al., 2009). The used SLR methodology has four phases (Khorasani et al., 2024), which are elaborated in the following.
3.1. Specifying the scope and research questions
Research questions (RQs) are the foundation of a systematic study. In fact, an SLR is an effort to answer RQs. In this study, we have defined 5 research questions as follows:
RQ1: How active is the field of WaaS brokers? This question has been answered in Section 3.4.
RQ2: What is the taxonomy of designing a WaaS broker, and how can related work be classified in this taxonomy? This question has been answered in Section 4.
RQ3: Which techniques have been used for solving the MWSP and to what extent? This question has been answered in Section 4.2.7.
RQ4: Where are the research articles published? This question has been answered in Section 3.4.
RQ5: What are the open issues in the field of WaaS brokers? This question has been answered in Section 5.
Our review covers a period of 20 years, from 2005, the advent of cloud computing, to the end of 2024. We have considered research works that explicitly addressed the design of a WaaS platform for cloud computing environments, or proposed a multiple workflow scheduling algorithm from the client-side perspective in cloud environments.
3.2. Planning the search process
The first step is specifying search strategies. In this review, we used a three-tier search strategy: database search, backward snowballing, and forward snowballing. Database search is a common method in literature surveys, where the researcher searches well-known online databases using search strings. To conduct the database search, we used IEEE Xplore, ScienceDirect, SpringerLink, ACM Digital Library, and Wiley Online Library. After reviewing some relevant papers and through a trial-and-error procedure by the authors, the following three queries were used as search strings:
Q1: (”multi-workflow” OR ”multi Workflow” OR ”multiple workflow” OR ”workflow ensemble”) AND cloud AND scheduling
Q2: workflow AND broker AND cloud AND scheduling
Q3: (”workflow as a service” OR ”WfaaS” OR ”WaaS” OR ”workflow-as-a-service”) AND cloud
These search strings were applied to the Title, Abstract, and Keywords sections of the papers to retrieve all relevant papers. All extracted papers and their venues (i.e., journals or conferences) were evaluated against predefined criteria (see the next sub-section), and the papers meeting those criteria were added to the set of included papers, which was then sent to the snowballing step.
Snowballing is done in two steps: backward and forward. Backward snowballing involves searching the reference list of an article to find new related articles. Forward snowballing involves searching among the citing papers of an article to find new related articles. To find the citing papers for an article, we used Google Scholar. The newly extracted articles underwent the evaluation process, and those added to the included papers were sent back to the snowballing step, continuing until no new articles were added to the set of included papers.
3.3. Planning the study selection process
The first step of this phase is determining the criteria for venue inclusion/exclusion to ensure that high-quality venues are selected. For journals, those indexed in the Journal Citation Reports (JCR) (ranks Q1-Q4) are included. For conferences, those indexed in the ICORE conference ranking (Computing Research and Education Association of Australasia, [n. d.]) (ranks A, B, and C) are added to the list of included conferences. All extracted studies that are not published in the included venues are discarded from further consideration.
The next step is defining the study selection strategy. Initially, the relevance of a paper is evaluated by reviewing its title, abstract, and keywords. If these elements do not provide sufficient information to make a definitive decision about its relevance, the full text of the paper is reviewed. For a paper to be included, it should explicitly design a workflow as a service or propose a multiple workflow scheduling algorithm for the cloud environment. Papers proposing scheduling algorithms for a single workflow, or those designed for other distributed environments such as clusters, grids, edge, or fog, have been excluded. Furthermore, papers that propose scheduling algorithms from the cloud provider’s point of view, which schedule on physical machines (such as (Wang et al., 2023, 2024; Fan et al., 2023a)), are excluded. A few papers have also been excluded because, although they claimed to be designed for the cloud environment, they did not consider important properties of the cloud, such as elasticity and scalability.
3.4. Conducting the SLR and analyzing results
After establishing the strategies, we proceeded with the SLR by implementing these strategies. As a result, we identified 43 venues (31 journals and 12 conferences), along with 74 studies that were included, which are detailed in the provided Excel supplementary file. Figure 2 presents a brief overview of the studies included per year and per publisher. As seen in Figure 2(a), the number of published papers shows an increasing trend in recent years, reaching its peak in 2024. This suggests that researchers have paid more attention to this topic in recent years. Please note that the figure starts from 2014, as only two papers were published before in 2009 and 2011. Figure 2(b) illustrates that IEEE journals and conferences have the highest number of papers in this area, followed by Springer and Elsevier. The top three venues with the most published papers are The Journal of Supercomputing with 7, and Future Generation Computer Systems and IEEE Transactions on Services Computing both with 6 published papers.
(a) Per year
(b) Per publisher
Figure 2. Number of included papers
The final step involves analyzing and classifying the data to address the research questions of the SLR. The primary objectives of this SLR are to identify the taxonomy of the WaaS broker and to classify the included papers within this taxonomy. After a thorough review of the included papers, and based on the proposed architecture of the WaaS broker in Figure 1, a taxonomy was developed for each of the four components of a WaaS broker. All included papers were carefully analyzed by two authors and classified according to the taxonomy. Any disagreements were reviewed and resolved by all authors. The results of the analysis are presented in Section 4.
4. Taxonomy and survey
In this section, we elaborate on the proposed taxonomy and then survey the related work based on this taxonomy. The taxonomy has been designed around the different components of a WaaS broker, and we will discuss it in the following four sub-sections, each focusing on a specific component. To conserve space in this SLR, a list of all included articles along with their classification results has been provided as a supplementary file in the form of an Excel document. This also facilitates easier searching and review of the included articles for readers.
4.1. Negotiation and pricing taxonomy
{forest}
[…]
Figure 3. Negotiation and pricing taxonomy
Figure 3 presents the proposed taxonomy for the negotiation and pricing components, which is divided into four categories. Below, we provide a detailed explanation of each category.
4.1.1. QoS Requirements
Users have varying requirements regarding quality of service criteria (e.g., execution time, cost, privacy, reliability, etc.; see Section 4.2.5). These requirements can be classified into two categories. The first category is Multi-criteria, where the user seeks to optimize multiple criteria simultaneously. In this case, the broker faces a multi-criteria or multi-objective optimization problem, which can be addressed by finding Pareto Optimal solutions or by optimizing a weighted sum of the criteria. Many studies in this category focus on minimizing both the makespan and cost of workflow execution (Wang et al., 2019; Xia et al., 2023; Li et al., 2022a; Li et al., 2024d), while others consider additional objectives, such as makespan, cost, and energy at the same time (Tarafdar et al., 2023). For example, Sun et al. (Sun et al., 2023) addressed a tri-objective multi-workflow scheduling problem in a hybrid cloud, aiming to minimize the total tardiness (i.e., the maximum between zero and the makespan of the workflow minus its deadline), total energy consumption in the private cloud, and the total renting cost in the public cloud. Similarly, Li et al. (Li et al., 2024b) considered the optimization of three objectives simultaneously: makespan, cost, and fairness.
On the other hand, the Constrained class focuses on optimizing one criterion while treating the other criteria as constraints. Two common constraints in this class are deadline-constrained and budget-constrained requirements. In a deadline-constrained scenario, the broker aims to meet the user-defined deadline while optimizing another criterion, typically minimizing the cost of workflow execution (Taheri et al., 2023; Rodriguez and Buyya, 2018). In a budget-constrained scenario, the broker seeks to keep the execution cost within the user-defined budget while optimizing other criteria, usually minimizing the workflow execution time (Li et al., 2023b, 2024a). Some studies address both deadline and budget constraints simultaneously (Arabnejad and Barbosa, 2017a, b; Rezaeian et al., 2019). Finally, there are also studies that optimize multiple objectives while ensuring that certain constraints are met. For example, Karmakar et al. (Karmakar et al., 2024) optimized resource utilization and rental costs while ensuring that the workflows met their deadlines.
(a) QoS Requirements
(b) Workflow Arrival
Figure 4. Number of articles on QoS Requirements and Workflow Arrival.
4.1.2. Workflow Arrival
Workflow arrival patterns can be categorized into four classes. In the online (or random) arrival pattern, workflows arrive at random times, with no predefined schedule (Chen et al., 2021; Liu et al., 2021; Liu et al., 2024; Taghinezhad-Niar et al., 2022). As a result, the broker must be prepared to handle workflow arrivals at any time. In contrast, batch arrival occurs when a set of workflows arrives together and is sent to the scheduler to be executed simultaneously (Wu and Yuandou, 2018; Zhou et al., 2024; Li et al., 2024c). Some articles that claim to adopt an online arrival pattern actually aggregate incoming workflows within a time interval and send them to the scheduler as a batch. These articles are classified under the batch category in this review (Li et al., 2022d; Li et al., 2024d).
The third category is ensemble arrival, which is similar to the batch arrival pattern but with a key distinction: all workflows in an ensemble share the same structure but have different sizes and input data (Zolfaghari and Abrishami, 2022; Malawski et al., 2015; Chen et al., 2024b; Jiang et al., 2015). Typically, users specify QoS requirements, such as deadlines and budgets, for the entire ensemble rather than for individual workflows. The final category is periodic arrival, where multiple workflows are released at regular intervals, with each having a defined period between consecutive instances. Generally, an instance of a workflow must complete before the next one is released, with the release of the next instance acting as a deadline for the previous one. Only two of the reviewed articles address this arrival pattern (Taheri et al., 2023; Chen et al., 2020a), and both consider using the reserved pricing model for running periodic workflows.
4.1.3. Negotiation
Negotiation is the process where the user and the WaaS broker agree on an SLA, such as pricing or execution time. For instance, in a deadline-constrained case, the user may specify a desired deadline for workflow completion, and the broker and user can then agree on a price. Negotiation can be divided into two types: static and dynamic. In static negotiation, the user specifies their QoS requirements, and the broker checks if it can fulfill them. If it can not, the request is declined, but if it can, the broker sends back a price for the user to accept or reject. However, standards like WS-Agreement Negotiation (Battré et al., 2010) enable the broker to negotiate QoS and price dynamically with users. For example, in a deadline-constrained scenario, if the broker cannot meet the specified deadline, it might propose a feasible deadline to the user. Or if the price is too high, the broker could offer a lower price in exchange for a more flexible deadline. Unfortunately, dynamic negotiation has not been employed by any of the reviewed articles, and only one article has implemented static negotiation. Thus, developing negotiation strategies could be a promising area for future research in WaaS broker design.
Among the reviewed studies, only Taghavi et al. (Taghavi et al., 2023) proposed a static negotiation module for a WaaS broker. They designed a deadline-constrained broker where the user submits a workflow with a specified deadline. The broker calculates the workflow’s critical path and accepts it if the deadline is sufficiently distant from the critical path (based on parameters set by the broker’s administrator). The broker then calculates the price for execution, considering factors like the deadline, resource availability, costs, and the broker’s profit margin. This price is presented to the user for approval or rejection. The negotiation component shows significant potential as a subject for future research directions.
4.1.4. Pricing
Pricing is the process of determining the cost of the services offered by the WaaS broker. To set the price, the broker must account for the actual cost of executing the workflow, its overhead costs (such as penalties for violating the SLA), its profit margin, and the user’s benefits. Typically, the price should be lower than the cost of the user executing the workflow directly by leasing resources from the cloud provider. When using on-demand instances, the broker is expected to reduce costs by reusing leased resources for multiple workflows. However, the broker can significantly lower execution costs by using reserved and spot instances. In this case, the pricing is determined based on broker-side factors, such as available idle reserved instances or the current spot instance price, along with client-side factors, such as the specified deadline for the workflow.
Pricing can be approached in two ways: pre-pricing and post-pricing. In pre-pricing, the broker sets the price for executing the workflow before it begins. This involves predicting the execution cost by examining the workflow’s structure, its required QoS, and the current availability and pricing of resources. Specifically, when using reserved or spot instances, the broker must assess the idle capacity of reserved instances and estimate the future cost of spot instances used for execution. The broker then uses this predicted cost to establish a final price, factoring in hidden costs and its profit. Pre-pricing is also employed during the negotiation phase to finalize the price with the user ahead of execution. In contrast, post-pricing involves determining the final price based on the actual cost incurred during workflow execution. While post-pricing offers greater accuracy and is more beneficial for the broker, users typically prefer knowing the price in advance to decide whether to accept or reject the offer.
Although the majority of the papers reviewed focus on the execution cost of workflows, either by minimizing the cost or staying within the user-defined budget, only one paper addresses the pricing policies of WaaS brokers. Taghavi et al. (Taghavi et al., 2023) introduce a post-pricing policy, called constant profit, and two pre-pricing policies: constant discount and prediction-based, for a deadline-constrained WaaS broker that utilizes a combination of on-demand and spot instances for workflow execution. In the constant profit model, after completing the workflow, the broker calculates the total execution cost and adds a fixed profit margin. In the constant discount approach, the broker estimates the workflow’s execution cost using on-demand instances and then applies a discount based on past experience with spot instances before presenting the final price to the user. The prediction-based pricing method involves the broker estimating the cost of execution ahead of time, using the current prices of on-demand instances and the anticipated cost of spot instances during the workflow’s runtime. These pricing strategies highlight a promising area for further development in WaaS brokers.
4.2. Task scheduling taxonomy
{forest}
[…]
Figure 5. Task scheduler taxonomy.
Figure 5 illustrates the proposed taxonomy for task scheduler components, comprising seven categories. Each category is detailed in the following sections.
4.2.1. Task Type
The task type defines the degree of flexibility in the task’s execution time. Some studies consider tasks to be rigid, meaning they require a fixed amount of computational resources, with execution time remaining (nearly) constant on these resources (Taghavi et al., 2023; Wang et al., 2014). However, most studies focus on flexible tasks, which exhibit varying performance depending on the resources used. To represent flexibility, some studies assume the execution time of each task on each resource is pre-determined through techniques like profiling (Wang et al., 2019). Others assume that each task has a known execution time on the fastest resource, with other resources assigned a weight reflecting their relative computational power. Based on this weight, the task’s execution time on any resource can be derived (Taghinezhad-Niar and Taheri, 2023; Liu et al., 2024). Lastly, some studies estimate the task size in Million Instructions (MI) and the resource’s CPU capacity in MIPS (Million Instructions Per Second), calculating the task’s execution time by dividing these two values (Rodriguez and Buyya, 2018; Tarafdar et al., 2023).
4.2.2. Execution Time Estimation
Most proposed scheduling algorithms rely on an estimation of each task’s execution time. Many studies assume a fixed execution time (Arabnejad and Barbosa, 2017b; Esteves and Veiga, 2015), which is impractical in dynamic environments like the cloud, where resource performance can be unpredictable. To address this, some researchers adopt probabilistic execution times for workflow tasks, modeling task execution time as a stochastic variable with an independent normal distribution characterized by a specific mean and variance (Ma et al., 2021; Liu et al., 2021; Xu et al., 2023). For instance, Chet et al. (Chen et al., 2021) model both task execution time and data transfer time among tasks as random variables and use their quantiles for approximation. A few studies manage execution time uncertainty using fuzzy logic. Zhu et al. (Zhu et al., 2020) represent task execution times and data transfer times using triangular fuzzy numbers. Based on these representations, they calculate fuzzy start and completion times for each task. They then introduce a list-based heuristic algorithm, Fuzzy Task Scheduling (FTS), which uses fuzzy priorities to schedule tasks effectively.
4.2.3. Task Concurrency
Task concurrency defines the number of tasks that can be executed concurrently on the same resource. Most studies adopt a single mode, where only one task is scheduled on each resource at a time (Chen et al., 2021; Taghinezhad-Niar and Taheri, 2023). However, some studies explore the multiple mode, which permits the concurrent execution of several tasks on the same resource. For instance, in (Taghavi et al., 2023; Saeedizade and Ashtiani, 2021), brokers rent large virtual machines (VMs) with a high number of CPU cores and significant memory, bundling tasks together and scheduling them on these VMs as containers. Similarly, Silva et al. (Silva et al., 2024) utilize multi-threading on a single node to enable the simultaneous execution of multiple tasks.
4.2.4. Data Transmission
Data transmission outlines how output data from parent (predecessor) tasks is delivered to child (successor) tasks. In the direct (or peer-to-peer) approach, data is transferred straight from the resource executing the predecessor task to the resource assigned to the successor task. This method requires synchronous communication, meaning the successor task must already be scheduled, and its resource allocation must be determined beforehand. A notable advantage of this approach is that data transfer time between tasks scheduled on the same resource becomes negligible. However, it may not be ideal for scheduling strategies that delay the execution of successor tasks. In contrast, the indirect (or shared) method involves the predecessor task storing its output data in shared storage, such as Amazon S3 or Amazon EFS. The successor task can then retrieve this data asynchronously, enabling more flexibility in task scheduling. Researchers have utilized both direct (Chen et al., 2018; Arabnejad et al., 2019) and indirect (Rajasekar and Palanichamy, 2022, 2021) methods in their studies.
4.2.5. Scheduling Criteria
This sub-section explores the scheduling criteria frequently used as optimization objectives or constraints in the MWSP. The most common criteria are the execution time of workflows and the rental cost of resources, which are addressed in almost all studies, either as optimization goals or constraints (such as deadlines or budgets). In addition to these, some studies focus on maximizing resource utilization (Liu et al., 2024; Chen et al., 2016; Jiang et al., 2011). Another key objective is minimizing energy consumption, which is generally a provider-focused goal. Although it is rarely considered a client- or broker-side concern, some studies suggest that brokers could reduce energy usage by adjusting the CPU frequencies of leased resources (Taghinezhad-Niar and Taheri, 2023; Taghinezhad-Niar et al., 2022). In hybrid cloud environments, where both private and public clouds are used, researchers often aim to reduce energy consumption within the private cloud while disregarding energy usage in the public cloud, as it is not typically a user concern (Li et al., 2022d; Sun et al., 2023).
Fairness is another criterion considered in some studies, although its definition varies. Some works, such as (Chen et al., 2021; Wu and Yuandou, 2018), approach fairness as a resource-centric measure to ensure balanced workload distribution across resources. In contrast, others, including (Li et al., 2024b, c; Rezaeian et al., 2019), define fairness as a user-centric measure that focuses on equitable resource allocation among workflows. For example, (Li et al., 2024b) introduces the concept of a workflow’s loss rate, which combines two factors: slowdown (the ratio of a workflow’s completion time when scheduled alongside others to its completion time when scheduled alone) and overspending (the cost ratio under the same conditions). Fairness is then defined as minimizing the disparity in loss rates across all workflows. Additional objectives considered by researchers include improving reliability (Taghinezhad-Niar and Taheri, 2023; Alam et al., 2024; Sampaio and Barbosa, 2020), ensuring privacy (Sun et al., 2023; Sharif et al., 2014), and enhancing security (Alam et al., 2024; Roy et al., 2024).
4.2.6. Scheduling Type
Workflow scheduling can be classified into two primary types: static and dynamic (Rodriguez and Buyya, 2017). Static scheduling involves assigning tasks to resources prior to the start of workflow execution. This approach typically produces high-quality solutions since it considers crucial structural aspects of the workflow, such as the critical path. However, static scheduling is limited by its reliance on precise runtime estimations and its inability to adapt to unforeseen changes in resource availability, making it less suitable for the uncertain nature of cloud environments. On the other hand, dynamic scheduling makes decisions during runtime, assigning tasks as they become ready—that is, once all their predecessor tasks have been completed. This enables dynamic methods to respond effectively to changes and uncertainties in the cloud environment. Nevertheless, these algorithms often make greedy, task-specific decisions, lacking a global perspective on the entire workflow. Consequently, the solutions generated by dynamic approaches are generally less optimal compared to those of static methods. To address the shortcomings of both approaches, hybrid strategies have been proposed. These methods begin by creating a static execution plan for the workflow and then apply dynamic adjustments during runtime to account for environmental uncertainties. This combination leverages the strengths of both static and dynamic scheduling to produce more robust solutions.
In static multiple workflow scheduling, two primary approaches have been used: one-by-one and combined. In the one-by-one method, workflows are scheduled sequentially, based on priority, while considering resource availability. For example, Rezaeian et al. (Rezaeian et al., 2019) apply the one-by-one approach using four different workflow selection methods: Ordered, which schedules workflows based on a predefined priority; round-robin, which iterates over workflows, scheduling one task from each workflow at a time; weighted round-robin, where workflows are selected according to a weighted formula before scheduling a task; and unfairness heuristic, which selects workflows based on their level of unfairness. Rimal and Maier (Rimal and Maier, 2017) employ three different workflow selection strategies: FCFS (First Come First Serve), EASY Backfilling, and MCT (Minimum Completion Time). In contrast, the combined approach merges all workflows into a single large workflow before scheduling it on available resources. References (Alam et al., 2024; Ahmad et al., 2022) use a similar approach, where they break down each workflow into levels based on task depth. Then, they create partitions where each partition contains tasks at the same level across different workflows. Other studies merge workflows by adding a dummy entry task to connect the input workflows’ actual entry tasks, and a dummy exit task to connect the workflows’ exit tasks (Taheri et al., 2023; Qin and Shao, 2025; Zheng et al., 2019).
Two strategies are commonly used for dynamic multiple workflow scheduling: immediate and periodic. In the immediate strategy, tasks are scheduled as soon as they become ready. However, this approach faces two challenges: first, the runtime overhead caused by executing the scheduling algorithm for each task as it becomes ready, and second, the greedy decision-making process for each task without considering the future tasks and their potential impact on the current scheduling. To address these issues, the periodic strategy was introduced. This method divides time into intervals, collecting all ready tasks within each interval and scheduling them all at once at the end of the interval. While this approach resolves the problems associated with immediate scheduling, it also introduces a delay, causing ready tasks to wait until the end of the interval, potentially missing deadlines. Most previous work in dynamic scheduling adopts the immediate strategy (Liu et al., 2021; Ma et al., 2021; Rajasekar and Palanichamy, 2022), although some utilize the periodic strategy (Taghavi et al., 2023; Lin et al., 2024). Chen et al. (Chen et al., 2017) introduced a hybrid strategy, where they use periodic scheduling but select tasks at the start of each interval that are predicted to have their earliest start time within the current interval. These tasks are scheduled on available resources ahead of time.
Only two studies have employed a combination of both static and dynamic scheduling types. Taheri et al. (Taheri et al., 2023) proposed a framework consisting of two phases: static planning and dynamic scheduling. In the static planning phase, a heuristic algorithm is used to schedule a combined workflow. During the dynamic scheduling phase, the system seeks to provision the necessary resources and execute tasks based on the static plan. However, adjustments may be made to the initial schedule to account for uncertainties in the cloud environment. Similarly, Zhou et al. (Zhou et al., 2017) introduced a comparable framework with an offline optimization stage using breadth-first search or A* algorithms to create a static schedule, followed by an online adaptation stage that refines the primary schedule using runtime information about cloud resources and workflow executions. Figure 6(a) illustrates the distribution of articles across each category. It is evident that immediate dynamic scheduling is the most commonly used approach.
(a) Scheduling types
(b) Scheduling algorithms
Figure 6. Number of articles for different scheduling types and algorithms
4.2.7. Scheduling Algorithm
The MWSP in a cloud environment is a well-known NP-hard problem (Ullman, 1975). Various algorithms have been employed by researchers to address this issue, with heuristic methods being the most commonly applied. The term heuristic refers to a method based on experience or judgment, designed to provide a solution that is close to optimal for a given problem (Silver, 2004). The majority of these studies have utilized the list heuristic to solve the MWSP (Chen et al., 2021; Zhou et al., 2018; Wang et al., 2021; Arabnejad et al., 2019; Xu et al., 2009; Ye et al., 2023). This heuristic involves two stages: task selection and resource mapping. In the task selection phase, each workflow task is assigned a rank according to a specific formula, after which the tasks are sorted and selected one by one based on their ranks. In the resource mapping phase, each selected task is assigned to a resource that optimizes the scheduling objective. Another commonly used heuristic in this field is the clustering heuristic (Li et al., 2024b; Ahmad et al., 2022), where tasks of the workflow are grouped into clusters, and each cluster is scheduled on a single resource to minimize data transmission time between tasks (in direct transmission mode). Other heuristic algorithms, such as A*, have also been explored by researchers (Zhou et al., 2016; Zhou et al., 2017).
Meta-heuristic methods are another popular class of algorithms for task scheduling. Meta-heuristics are higher-level heuristics that are problem-independent and provide generic frameworks for finding near-optimal solutions to optimization problems (Abdel-Basset et al., 2018). Li et al. (Li et al., 2022c) proposed the Multi-swarm Co-evolution-based Hybrid Intelligent Optimization (MCHO) algorithm, which aims to minimize makespan and cost while meeting the deadline of each workflow. They utilized the Particle Swarm Optimization (PSO) algorithm, incorporating three different swarms: the first and second swarms focus on optimizing makespan and cost, respectively, while the third swarm is introduced to balance the two objectives. Additionally, they employed Genetic Algorithms (GA) as an elite enhancement strategy for each swarm, aiming to exploit elite individuals and increase the likelihood of finding more non-dominated solutions. Non-dominated Sorting Genetic Algorithm III (NSGA-III) is also applied to address the MWSP as a prominent multi-objective meta-heuristic method (Li et al., 2024b, c). Li et al. (Li et al., 2024b) first introduced a clustering heuristic called Depth-First-Search considering Critical Subsequent Task (DFS-CST) to partition each workflow into clusters and then applied NSGA-III to assign resources to each cluster, aiming to minimize three objectives simultaneously: execution time, execution cost, and unfairness.
Zhou et al. (Zhou et al., 2024) introduced the Evolutionary Multitasking Constrained Optimization-Adaptive Double-level knowledge Transfer (EMCO-ADT) method for deadline-constrained multi-workflow scheduling. They treat each workflow scheduling problem as an individual ”task” that can be addressed using Evolutionary Algorithms (EAs). To enhance the performance of multi-workflow scheduling, they incorporate cross-task knowledge transfer between independent workflow schedules. This approach utilizes the Evolutionary Multi-Task Optimization (EMTO) paradigm, which facilitates knowledge extraction and transfer across distinct tasks. Qin and Shao (Qin and Shao, 2025) proposed a bi-population co-evolutionary algorithm for multi-objective, deadline-constrained multi-workflow scheduling. Their method involves two sub-populations: a deadline-aware sub-population that focuses on finding feasible schedules that meet deadlines, and a multi-objective sub-population that aims to minimize both makespan and cost. After each iteration, a bi-directional exchange of schedules occurs between the two sub-populations, promoting the diversity of each population.
Other meta-heuristic algorithms, such as the Owl Search Algorithm (OSA) (Li et al., 2022d), Poor and Rich Optimization (PRO) algorithm (Li et al., 2024d), the Snake Optimizer (SO) (Li et al., 2023b), and the Gradient-Based Optimizer (GBO) (Li et al., 2024a), have also been applied to address the MWSP. Additionally, three studies have utilized Genetic Programming Hyper-Heuristic (GPHH) to generate heuristic rules for multi-workflow scheduling (Fan et al., 2023b; Sun et al., 2024; Escott et al., 2022). In this approach, each heuristic is represented as a syntax tree (GP tree), where the leaves are terminal nodes corresponding to problem parameters such as task execution time or a VM’s execution speed or cost, and the internal nodes represent function nodes, such as addition or maximum. Each heuristic is treated as an individual evolved by GPHH, and ultimately, the best-performing heuristic(s) is selected for scheduling.
Recently, Reinforcement Learning (RL) has been applied to solve the MWSP. RL is a machine learning approach where an agent learns through interaction with an environment, making decisions based on feedback from its actions and experiences. The agent selects actions based on its policy and the environment’s state, receives rewards or penalties as feedback, and adjusts its policy accordingly. Li et al. (Li et al., 2022a) use a Double Deep Q-Network (DDQN) to minimize both makespan and cost in a multi-workflow scheduling context. They employ a two-level scheduling strategy: at the first level, a ready task is selected for scheduling, and at the second level, a suitable VM is chosen for that task. Each level features two sub-agents: a time sub-agent and a cost sub-agent, aimed at optimizing time and cost, respectively. After training, the model achieves an optimal balance between makespan and cost after 400 episodes. Pan and Wei (Pan and Wei, 2024) proposed a similar dynamic scheduling algorithm using a Deep Q-Network (DQN), but focusing solely on the VM selection phase. In another related study, Huang et al. (Huang et al., 2022) utilize a Deep Neural Network (DNN) for VM selection of the ready tasks, combined with an evolutionary algorithm to optimize the DNN’s parameters.
Lin et al. (Lin et al., 2024) also employed RL for workflow task scheduling across a combination of on-demand and spot instances, though in a different way. For task scheduling, they used a DNN that takes an 11-dimensional vector of the ready task’s features as input. Instead of selecting a specific VM for task execution, the output is a meta-policy. The meta-policies are as follows: Util-priority: schedule the task on the available spot instance with the least available resources; Avl-priority: schedule the task on newer spot instances; To-spot: launch a new spot instance and schedule the task on it; To-on-demand: schedule the task on an idle available on-demand instance, or if none are available, launch a new on-demand instance; Wait: defer the task’s execution until the next scheduling period. While all the aforementioned works proposed dynamic scheduling algorithms, Zhang et al. (Zhang et al., 2023) introduced a static scheduling algorithm for multi-workflows. Their approach first utilizes a GA meta-heuristic to determine the optimal ordering of the workflow’s tasks, and then employs a DQN agent to assign tasks to VMs.
Multi-criteria decision-making (MCDM) is a field that focuses on choosing the best alternative from several options by evaluating multiple criteria or attributes (Taherdoost and Madanchian, 2023). Among the popular MCDM methods, only the Technique for Order of Preference by Similarity to Ideal Solution (TOPSIS) has been applied to the MWSP (Xia et al., 2023; Chakravarthi and Shyamala, 2021; Chakravarthi et al., 2022). Chakravarthi et al. (Chakravarthi and Shyamala, 2021; Chakravarthi et al., 2022) use TOPSIS to determine the best VM for a ready task in a dynamic scheduling algorithm. They input a $m\times r$ decision matrix into TOPSIS, where $m$ represents the number of available VMs, and $r$ is the number of criteria (3 in this case): execution time, transfer time, and execution cost. Another approach is Game theory, which is used to model scenarios where multiple independent players interact and influence each other’s outcomes (Nisan et al., 2007). Among the reviewed studies, only one uses this method. Wu and Yuandou (Wu and Yuandou, 2018) apply multi-stage dynamic game theory to minimize makespan while enhancing fairness and utilization. Each optimization objective—makespan, fairness, and utilization—is considered as a competing player. The Nash equilibrium is determined as the optimal solution to the problem.
Finally, mathematical optimization (or mathematical programming) utilizes techniques like Linear/Nonlinear Programming (LP/NLP) to address optimization problems. Karmakar et al. (Karmakar et al., 2024) apply NLP to minimize execution costs for a set of deadline-constrained workflows. They first identify the critical paths within the workflow and then aim to minimize the total required MIPS for tasks on these critical paths, ensuring the workflow’s deadline is met. The problem is formulated as an NLP, and an approximate solution is found using the Lagrange Multiplier Method. Figure 6(b) shows the distribution of articles for each algorithm. It is clear that heuristic algorithms are the most frequently used, followed by meta-heuristic algorithms.
4.3. Resource provisioner taxonomy
{forest}
[…]
Figure 7. Resource provisioner taxonomy.
Figure 7 depicts the proposed taxonomy for the resource provisioner component, which is organized into six categories. A detailed explanation of each category follows.
4.3.1. Resource Type
As shown in Figure 8(a), the majority of the articles reviewed schedule workflows directly on VMs, which are traditional IaaS resources (Chen et al., 2020b). However, containers have recently emerged as a lightweight virtualization technology, allowing users to bundle tasks as standalone, isolated applications. Consequently, some articles have adopted containers for task scheduling on VMs. For example, in (Taghavi et al., 2023; Li et al., 2022c), the broker rents large VMs with numerous CPU cores and ample memory, using containers to bundle and execute multiple tasks on a single VM. Although serverless computing and FaaS have emerged as promising models for executing scientific workflows in the cloud (Malawski et al., 2020), only Sun et al. (Sun et al., 2024) have explored the scheduling of multiple workflows using a combination of serverless functions and VMs. This combination shows promise and will be discussed further in future directions (Section 5).
(a) Resource Type
(b) Resource Pricing
Figure 8. Number of articles for different resource types and pricing
4.3.2. Resource Diversity
Most of the papers reviewed focused on a cloud provider offering heterogeneous resources with varying configurations (such as CPU cores, memory, and storage), aiming to identify the optimal resource configuration for each task based on its requirements and scheduling objectives and constraints (e.g., deadlines and budgets). However, some researchers opted for a homogeneous environment, using a single resource type for all tasks, typically to simplify the scheduling process and reduce the search space (Liu et al., 2024; Zhu et al., 2020). To justify this approach, Malawski et al. (Malawski et al., 2015) argue that, for most applications, there will generally be one or two VM types that offer the best price/performance ratio. In contrast, Taheri et al. (Taheri et al., 2023) utilized only a single type of large VMs, then employed containers to schedule tasks on these VMs.
4.3.3. Resource Pricing
As depicted in Figure 8(b), the majority of the reviewed articles utilized the widely known on-demand pricing model for executing users’ workflows, leveraging the elasticity of cloud resources (Adhikari and Koley, 2018; Shen et al., 2025). However, some studies, either implicitly or explicitly, considered a fixed pool of reserved resources, which doesn’t fully capture the essence of cloud computing (Pan and Wei, 2024; Zhang et al., 2023). Despite this, Chen et al. (Chen et al., 2020a) used reserved resources for executing periodic workflows, proposing a heuristic algorithm to determine the optimal amount of reserved resources. More interestingly, several studies explored how the WaaS broker can reduce execution costs by combining spot and on-demand instances (Taghavi et al., 2023; Zolfaghari and Abrishami, 2022; Sampaio and Barbosa, 2020; Chen et al., 2024b; Zhou et al., 2016), or reserved and on-demand instances (Taheri et al., 2023; Zhu et al., 2020).
Zolfaghari et al. (Zolfaghari and Abrishami, 2022) developed a multi-class workflow scheduling algorithm, where tasks are assigned to different resource classes based on their priority. Higher-priority tasks are sent to higher-class resources, which are reliable but expensive on-demand instances, while lower-priority tasks are allocated to cheaper spot instances with lower bid prices. If a task fails to execute within a specified time, it is moved to a higher-class resource until completion. Similarly, Zhou et al. (Zhou et al., 2016) combined spot and on-demand resources for workflow execution. They initially used the A* algorithm to identify the optimal on-demand instance type for each task to minimize costs while meeting the workflow’s deadline. Then, a refinement algorithm replaces some on-demand instances with spot instances to further reduce costs. Taheri et al. (Taheri et al., 2023) focused on scheduling periodic workflows using a mix of reserved and on-demand instances. Their scheduling algorithm prioritizes reserved instances for each task until its sub-deadline is reached. If resources are insufficient, the algorithm switches to on-demand instances to avoid missing the sub-deadline. Combining different pricing models to reduce execution costs is a promising strategy that will be discussed in future directions (Section 5).
4.3.4. Provision Timing
Provision timing is categorized into as-needed and periodic approaches. In as-needed provisioning, the broker acquires new resources when a task (or a group of tasks) requires them—typically because they cannot be scheduled on the currently available resources without violating an optimization criterion or constraint (e.g., a sub-deadline). In this approach, the resource provisioner is usually integrated within the task scheduler rather than being a separate module. As a result, the scheduling algorithm is responsible for both provisioning and de-provisioning resources. Whenever there is a resource shortage preventing the scheduling of a task, the algorithm triggers commands to lease new resources. The key advantage of this method is its prompt response to environmental events. For instance, in reference (Taheri et al., 2023), the scheduling algorithm calculates a parameter called laxity for each task, indicating the amount of time it can be delayed before missing its sub-deadline. If a task is ready to execute but no idle resources are available, the scheduler postpones it until the necessary resources become free, provided its laxity is still positive. If the required resources are not released and the laxity reaches zero, the scheduler proceeds to request new resources.
In contrast, periodic provisioning, commonly used in dynamic scheduling algorithms (whether immediate or periodic), requires the task scheduler to allocate workflow tasks to currently available resources. The resource provisioner operates as an independent module, periodically assessing the current resources based on parameters such as CPU load, task queue size, and other metrics. If necessary, the provisioner decides whether to lease additional resources or release some existing ones. One disadvantage of this method, compared to the as-needed approach, is that it responds to changes in the environment and workflow requirements later, typically at the end of each time interval. However, this delay can also be an advantage. In the as-needed approach, whenever a task requests a resource, it is immediately granted, often without considering the status of other tasks (either upcoming or current). This can result in prematurely leasing resources when, in a few moments, a suitable resource might become available, or it might be better to wait and request more powerful resources as new tasks arrive. By delaying resource provision until the end of each time interval, a more informed decision can be made, taking into account the overall status of resources and the tasks in the queue. However, this approach risks missing certain constraints, such as deadlines, so the broker must make more conservative decisions in these situations.
Taghavi et al. (Taghavi et al., 2023) combine dynamic periodic scheduling with periodic provisioning, where the provisioning intervals are significantly longer than the scheduling intervals (ranging from 5 to 10 times longer). During each provisioning interval, the provisioner checks whether the number of idle cores on the VMs is sufficient to accommodate the required cores for the ready tasks. If not, new VMs are leased from the cloud provider. In contrast, Malawski et al. (Malawski et al., 2015) implement dynamic immediate scheduling with periodic provisioning. Their approach computes resource utilization at the end of each interval and adjusts the number of leased VMs if the utilization exceeds or falls below certain thresholds. Tarafdar et al. (Tarafdar et al., 2023) employ a different approach to periodic provisioning, where, at the end of each time interval, they attempt to predict the number of workflows arriving in the next interval using historical data from previous intervals and scale the resources up or down accordingly.
4.3.5. Provisioning Delay
When a broker requests the provisioning of a VM from a cloud provider, it does not become available immediately. Instead, it takes a significant amount of time to be deployed on a physical host and for its operating system to boot. This period is referred to as provisioning delay. Even though containers, a lightweight virtualization technology, have a much shorter provisioning delay than VMs, they still experience some delay. This provisioning delay can be substantial, even in the case of containers, and thus, to make accurate provisioning and scheduling decisions, this parameter should be considered. In particular, provisioning delay plays a crucial role in provisioning algorithms, as in some situations it may be better not to release an idle resource so that it can be used for future tasks, avoiding the delay associated with provisioning new resources. To minimize VM provisioning delays, some researchers opt to use larger VMs and run multiple tasks on these VMs using container technology (Taghavi et al., 2023). However, only about 32% of the reviewed articles considered provisioning delay (Partheeban and Kavitha, 2019; Li et al., 2022d), which we refer to as delay-aware, while the remainder of the articles did not mention it at all (Ye et al., 2023; Li et al., 2024a), and are categorized as instantaneous.
4.3.6. Deployment Model
This taxonomy examines how the broker acquires its required resources, either from a single source or multiple sources. The majority of articles (66 papers) utilize a single-cloud provider to lease their resources. However, relying on a single provider has its drawbacks, including vendor lock-in, limited resource diversity, restricted pricing models, provider unavailability, and geographical location constraints (Khorasani et al., 2024). As a result, some studies (7 papers) adopt the multi-cloud model, where the broker leases resources from multiple, independent cloud providers (Taghinezhad-Niar and Taheri, 2023; Xu et al., 2023; Chen et al., 2020b; Sun et al., 2024). In contrast, a few articles (3 papers) implement a hybrid-cloud approach, where the broker uses resources from both a local private cloud and a public cloud, turning to the latter when local resources are insufficient (Li et al., 2022d; Sun et al., 2023; Sharif et al., 2014). References (Sun et al., 2023; Sharif et al., 2014) specifically use the private cloud to meet privacy requirements, where certain workflow tasks must be executed on a private cloud due to the sensitive nature of the data involved.
4.4. Monitoring and fault tolerance taxonomy
{forest}
[…]
Figure 9. Monitoring and fault tolerance taxonomy.
Resource failures are an inevitable aspect of computer datacenters, making it crucial for WaaS brokers to monitor resource performance and respond appropriately when failures occur. However, this requirement has not been adequately addressed in many reviewed studies. This is especially important in the spot pricing model, where the unreliability of resources and the potential for service providers to reclaim them during task execution are concerns. Despite these challenges, only 11 articles, including all five that utilize spot instances, address fault tolerance.
Figure 9 outlines a taxonomy for monitoring and fault tolerance. The simplest method, re-execution, involves transferring tasks from a failed resource to a new one, but this results in the loss of all progress made up to that point (Malawski et al., 2015; Esteves and Veiga, 2015; Sampaio and Barbosa, 2020; Zhou et al., 2016). This can be costly for the WaaS broker, particularly for scientific workflows with long-running tasks and strict deadlines. To minimize this cost, some papers use checkpointing (Taghavi et al., 2023; Zhou et al., 2024; Zolfaghari and Abrishami, 2022; Xu et al., 2023), where snapshots of tasks or VMs are periodically saved. If a failure occurs, the task is reassigned to a new instance, and the most recent snapshot is restored. Additionally, some cloud providers like AWS offer a two-minute notice before stopping or terminating spot instances, which allows enough time to take a snapshot. However, the process of taking and restoring snapshots can be time-consuming.
The final method is replication, which involves running duplicate instances of each task across multiple VMs to enhance the reliability of task execution (Taghinezhad-Niar and Taheri, 2023). In the event of a failure, the task can continue running without disruption, as its duplicate instances remain operational on other resources. While this approach ensures task continuity without losing time, it adds additional costs for the broker due to the need for extra resources to run the duplicated tasks. Chen et al. (Chen et al., 2024b) combine replication and re-execution for workflows running on a mix of on-demand and spot instances. They schedule tasks on spot instances only if there is sufficient time for re-execution in case the resource is reclaimed by the provider. Otherwise, the task is scheduled on an on-demand instance. When tasks are assigned to spot instances, they attempt to identify idle time on on-demand instances to replicate the tasks where possible. If a spot instance is reclaimed, tasks running on it are checked for duplicates. Tasks with duplicates are ignored, while those without duplicates are re-executed on another instance.
5. Future directions
Although WaaS brokers and the MWSP have made significant progress, there is still considerable potential for improving scheduling and provisioning algorithms. Specifically, there are areas that have received less attention and could benefit from further development, as well as new technologies, such as serverless computing, that can be leveraged. In the following, we outline some promising approaches for enhancing WaaS brokers, organized by the key components of a WaaS broker.
5.1. Negotiation and Pricing
Negotiation and pricing have not yet been sufficiently explored and remain open research challenges. Researchers could develop new negotiation algorithms to facilitate the creation of SLAs between brokers and users.The only prior work on negotiation mechanisms (Taghavi et al., 2023) involves receiving the user’s workflow deadline and proposing a price, which the user can accept or reject. Similar approaches can be explored for other constraints. For instance, in a budget-constrained scenario, the negotiation component could receive the user’s maximum budget and provide an estimated execution time for the workflow, which could then form the basis for an SLA. In multi-criteria optimization scenarios, the broker might identify several Pareto-optimal solutions and offer them to the user for selection. For example, in balancing execution time and cost minimization, the broker could present a list of non-inferior $\langle time,cost\rangle$ pairs for the user to choose from. The broker could also offer various options for QoS criteria, each associated with different prices, allowing users to select higher-cost, higher-reliability options for more critical or sensitive workflows. Additionally, dynamic negotiation protocols could be explored to create more user-friendly agreements.
Pricing policies also remain underexplored, presenting significant opportunities for further research. Particularly when the broker utilizes reserved or spot instances, pricing policies should incentivize users to submit workflows during off-peak times, when reserved instances are idle and spot instances are more affordable. Policies could also encourage users to set more flexible deadlines, allowing the broker’s scheduler to make better use of these cheaper resources. Furthermore, existing work in this area primarily employs cost-based pricing, where the total workflow execution cost is used to determine the price. Future research could explore more advanced pricing strategies, such as value-based or market-based pricing models (Wu et al., 2019), which are prevalent in economics and cloud computing, as potential foundations for more innovative pricing policies.
5.2. Task Scheduling
In the case of scheduling algorithms, while heuristic and meta-heuristic approaches have reached a certain level of maturity, other methods still hold significant potential for creating more effective scheduling algorithms. Game theory has been employed in only one previous study, yet it seems to have untapped potential for both scheduling and pricing algorithms. Additionally, mechanism design, a branch of game theory focused on designing rules that incentivize agents to act in ways that align with the system’s global objectives, has been successfully applied in scheduling multiple functions (Moakhar and Abrishami, 2024) and workflows (Mahdizadeh and Abrishami, 2024) in edge computing environments. This method could be a promising approach for tackling the MWSP. Furthermore, RL has recently gained significant attention in the scientific community. Although seven previous studies have already applied RL to solve the MWSP, it still offers great potential to propose better solutions to this problem.
Regarding task concurrency, only eight studies have explored running multiple tasks on a single resource. For a WaaS broker tasked with executing tens or even hundreds of tasks concurrently, it is impractical to rent a separate resource for each task. Therefore, the use of container technology to run multiple tasks on a single VM warrants more attention from researchers. Finally, for scheduling types, combining static planning with dynamic scheduling appears to be a very promising approach, as it leverages the strengths of both scheduling strategies. However, only two previous works (Taheri et al., 2023; Zhou et al., 2017) have employed this combination. Other researchers could explore this approach to develop more effective solutions.
5.3. Resource Provisioning
Since the WaaS broker aims to maximize its profits, it is logical to adopt various resource pricing models to minimize its costs. One effective approach is the combination of inexpensive but unreliable spot instances with more costly yet reliable on-demand instances, a strategy employed by six previous studies. However, there is still room for the development of new brokers that leverage this combination. Another promising option is the use of cheaper but non-elastic reserved instances in conjunction with elastic but more expensive on-demand resources. Surprisingly, only two prior works have explored this promising combination, and it is expected that more researchers will consider such models in the future. Additionally, combining all three pricing models—spot, reserved, and on-demand—has the potential to reduce costs while maintaining acceptable levels of reliability, yet none of the existing works have considered this comprehensive combination.
The emergence of serverless computing and the FaaS model offers new opportunities for WaaS brokers (Wen et al., 2023). This model eliminates the need for users and brokers to manage resources and handle scheduling. Instead, the cloud provider takes care of provisioning, scaling, and function scheduling. Additionally, FaaS operates on a utilization-based, fine-grained pricing system, charging users only for the actual resources they consume, which resolves concerns about idle resources. Although the application of FaaS for scientific workflows has been explored by some researchers (Malawski et al., 2020), the use of serverless platforms in workflow execution is still in its early stages. Most existing research focuses on frameworks designed to run a single workflow in a serverless environment. Notable examples include StepConf (Wen et al., 2022), Wukong (Carver et al., 2020), SWEEP (John et al., 2019), and StarShip (Basu Roy and Tiwari, 2024). In addition, several studies have investigated scheduling workflows in a multi-cloud environment, utilizing serverless services from various providers, such as AWS Lambda, Google Cloud Functions, and IBM Cloud Functions. Examples of such studies include xAFCL (Ristov et al., 2023), XFaaS (Khochare et al., 2023), SwitchFlow (Chen et al., 2024a), and FaaSt (Ristov and Gritsch, 2022).
However, there are some inherent characteristics of serverless computing that make running workflows difficult or inefficient (Roy et al., 2022). One key feature is the separation of computation and storage, which is designed to support auto-scaling for fluctuating workloads. As a result, serverless platforms use stateless functions that do not retain temporary state information in persistent storage, which creates challenges for workflows that require data storage and exchange between tasks. To address this, brokers must separately provision storage services such as AWS S3 or EFS for data exchange, but these options are slower and less efficient than direct data transfers between tasks (Li et al., 2023a). Additionally, serverless platforms are optimized for lightweight, short-running tasks, and typically impose task execution limits (e.g., 300 seconds for AWS Lambda and 60 seconds for Google Cloud Functions). This makes it difficult to run long-running, high-performance scientific workflows. Furthermore, serverless platforms use a scale to zero approach, where function containers are de-provisioned after a period of inactivity (e.g., 15 minutes in AWS). The cold start issue arises when a new function instance must be launched to handle an incoming request, resulting in delays. This cold start delay can be particularly problematic for short-running tasks, making execution inefficient.
These challenges have led researchers to consider hybrid execution strategies, combining serverful and serverless platforms to run scientific workflows (Malawski et al., 2020). In these hybrid platforms, the issues faced by serverful platforms, such as underutilization of leased VMs, are mitigated by the use of serverless platforms, and vice versa. In hybrid approaches, workflows are executed on a combination of traditional serverful (using VMs) and modern serverless (using functions) cloud platforms. VMs are well-suited for long-running or sequential tasks, while functions are more appropriate for lightweight, parallel tasks. Additionally, the local storage of the VM can be leveraged to share large datasets between tasks. When there is a significant volume of data to be transferred between two tasks, scheduling them on the same VM can eliminate data transfer delays, as they can use shared storage. Serverless services can be paired with any serverful pricing model (on-demand, spot, or reserved) to reduce execution costs and makespan. Although a few researchers have considered scheduling scientific workflows on hybrid platforms, most of them focus on scheduling a single workflow.
Roy et al. (Roy et al., 2022) proposed Mashup, a framework for executing scientific workflows in a hybrid environment that combines VMs and functions. They formulated an optimization problem to determine the optimal execution environment for each task in a workflow, aiming to minimize the overall execution time. The solution involved executing each task on its preferred platform (either a function or a VM). Zhang et al. (Zhang et al., 2024) presented the Budget-Constrained Workflow Scheduling (BCWS) algorithm, a heuristic approach for scheduling a single workflow with a limited budget on a hybrid setup of on-demand VMs and serverless functions, aiming to minimize makespan. The algorithm follows three phases. In the first phase, Initial Schedule Construction, the cheapest functions are allocated to all tasks, and an initial critical path is computed. If any budget remains, faster functions are assigned to critical path tasks to reduce the makespan, continuing until the budget is exhausted. In the second phase, Resource Replacement, the algorithm attempts to replace the rented functions with cheaper but non-slower VMs to restore the remaining budget. The third phase, Budget Reallocation, seeks to assign faster functions using a similar approach to the first phase, further decreasing the makespan.
Other platforms such as Lithops (Eizaguirre et al., 2024), DEWE v3 (Jiang et al., 2017), and HyperFlow on AWS Lambda and ECS (Elastic Container Service) (Balis et al., 2020) have also explored executing workflows in a hybrid environment combining serverful and serverless platforms. However, these platforms focus solely on executing a single workflow, which, while promising, is insufficient for developing a WaaS broker. The only prior work that addresses the execution of multiple workflows in a hybrid environment is Sun et al. (Sun et al., 2024), which is included in this SLR. The integration of serverless FaaS platforms with various pricing models (reserved, spot, and on-demand) for serverful VMs or containers presents a highly promising approach for provisioning resources, aiming to reduce costs and enhance performance. As a result, new provisioning and scheduling algorithms are needed to effectively leverage these emerging serverless platforms for scientific workflow execution.
5.4. Monitoring and Fault Tolerance
The Monitoring and Fault Tolerance component has also received limited attention from researchers. A combination of various fault tolerance methods can be applied in different scenarios. For instance, in a broker using reserved instances, the replication method can be employed when there are a high number of idle reserved instances. Alternatively, the re-execution method can be used for short-running tasks, while checkpointing can be applied for long-running tasks. Additionally, this component could work in conjunction with the negotiation and pricing component to offer different options to the user. For example, users could opt for a more reliable service for executing critical workflows by paying a higher fee. In such a scenario, the pricing component could calculate the cost of task replication, with assistance from the fault tolerance component, and incorporate it into the overall pricing.
6. Conclusion
This study conducts a Systematic Literature Review (SLR) on Workflow as a Service (WaaS) brokers in cloud environments, covering the period from 2009 to 2024. The SLR identified 74 relevant articles and 43 venues using a comprehensive search strategy. Based on the analysis of the extracted articles, a taxonomy has been proposed by considering the key components of WaaS brokers, namely negotiation and pricing, task scheduler, resource provisioner, and monitoring and fault tolerance. All articles have been classified and reviewed according to this taxonomy, and the findings are available in a supplementary Excel file.
The results reveal that WaaS brokers have garnered considerable attention from researchers, especially in the past three years. While significant progress has been made in the provisioning and scheduling components, other components, such as negotiation and pricing, remain underexplored, offering ample opportunities for improvement. Nonetheless, further advancements are possible even within the scheduling and provisioning domains. For instance, scheduling approaches like game theory and reinforcement learning hold promising potential for further refinement and application. In the provisioning domain, the combination of diverse pricing models (on-demand, reserved, and spot) remains insufficiently studied, presenting an opportunity for the development of novel models and algorithms. Moreover, emerging serverless service models, particularly Function as a Service (FaaS), and their integration with traditional serverful models such as VMs and containers, have introduced new avenues for WaaS broker research. These advancements create exciting prospects for researchers to explore innovative solutions and expand the capabilities of WaaS brokers in the evolving landscape of cloud computing.